OpenStudy (anonymous):

PLEASE PLEASE HELP AGAIN :( let P1 (-3, 4) and P2 (9, 10) be given points and suppose segment P1P2 is divided into 10 equal segments. determine the equation of a line perpendicular to P1P2 and passes through the point Q where W is a point on P1P2 six units from P2 i get how it's supposed to look like... but i'm having trouble finding Q distance of p1 and p2 is 6sqrt5 slope of p1 and p2 is 1/2 tried doing midpoints and all but still no success :(

OpenStudy (anonymous):

Well, let's see. P1 (-3,4) P2 (9,10) now, from the question, we know that Q is 4 units away from P1 and 6 units away from P2. Therefore, x must have moved \[\frac{4}{10}*12=\frac{24}{5}\] and y moved \[\frac{4}{10}*6=\frac{12}{5}\] from P1. So Q is \[( \frac{9}{5},\frac{32}{5})\] Now that you know the slope, the rest won't be that hard.

OpenStudy (phi):

Here is a picture  If you call the entire distance from p1 to p2 10 then you want to go 4/10 of the way. You know the x distance from p1 to p2 is 12 (9 - (-3))= 12 so set up a ratio to find how far to move in the x direction. you can do the same to find the y value: \[ \frac{y}{6}= \frac{4}{10} \]